# Chemical enrichment of metal-poor stars orbiting massive black hole companions

### v2

@@ 5.2 Altering the birth metallicity of main-sequence companions to LIGO-like black holes
Here we relax the constraints derived from the specific configuration of the BH3 system and expand our calculations to consider metal pollution in a $0.8\,M_{\odot}$ main-sequence star in a binary system with varying separations ($a$ from $0.01$ to $1R_{a}$) and black hole companion masses ($M_{\rm BH}$ from $10$ to $100\,M_{\odot}$). We present the results from this exercise in Figure 5, for which we have derived the total amount of iron accreted using the median value calculated from the particle trajectories of halo stars (S. Shen et al., 2017) for redshifts where $a/R_{a}<1$. We give, for simplicity, an initial birth [Fe/H] value of 5-5 to all stellar companions.
@@
This exercise allows us to probe a diverse collection of binary black hole-star systems that can inform future detections in the Milky Way. The range of stellar black hole masses in the Galaxy is closely linked to the life cycle of massive stars (e.g., F. Özel et al., 2010), as well as the energy and dynamics associated with supernova explosions (e.g., T. Sukhbold et al., 2016; S. L. Schrøder et al., 2018; A. Vigna-Gómez et al., 2024). By studying this distribution through observations (e.g., A. Olejak et al., 2020), we can gain insights into the types of black hole companions we might expect to find alongside low-mass stars. In the top panel of Figure 5, we plot the mass distribution of the LIGO-Virgo-Kagra black holes (R. Abbott et al., 2023b) and the black hole components of low mass X-ray binaries (F. Fortin et al., 2024). This distribution takes into account the masses of the primary and secondary components of the mergers, while black holes resulting from the merger were not considered. The low mass X-ray binary distribution peaks at around $15\,M_{\odot}$ while the LIGO distribution peaks at a larger value around $35\,M_{\odot}$.
@@
A few important points stand out from a close examination of Figure 5. For binaries hosting lighter black hole companions at large separations, the resulting surface abundances are low and closest to the star’s birth metallicity. For binaries hosting high-mass black hole companions at short separations, the surface abundance resulting from metal accretion can be significant. On the rightmost panel of Figure 5 we plot, for comparison, the metallicity distribution of the thick disc and halo stars from Gaia DR3 (B. Kim & S. Lépine, 2021). This bimodal distribution has a peak near 0.6-0.6, which is attributed to the thick disk stars, and a peak around 1.8-1.8 attributed to the halo stars. We thus conclude that accretion onto low-mass, main-sequence companions, when enhanced by a massive black hole companion, can play a prominent role in altering their surface abundances. In these cases, we expect the inferred surface abundances to trend with stellar luminosity, as mixing processes become progressively more dominant as the star evolves.
@@
Figure 6: The surface abundance enhancement via metal accretion of $0.8\,M_{\odot}$ companion stars with a given birth metallicity. The diagonal gray line shows the values for which the surface abundance is that of the star at birth. The color gradient indicates the different radii of the star as it evolves off the main sequence, with darker colors corresponding to more compact main sequence stars and lighter colors denoting extended giants. The orange curve denotes the surface abundance of stars in the main sequence. The horizontal teal dashed line shows the observed [Fe/H] value for Gaia BH3, and the teal star marker matches the value of the observed metallicity to its current radii. The horizontal green line shows the minimum value for [Fe/H] if at birth the star was solely made up of hydrogen and all of its iron content came from accretion from the ISM. The dotted vertical lines are critical metallicity for population II formation as derived by V. Bromm & A. Loeb (2003) and R. Schneider et al. (2006).
@@ 5.3 On the birth metallicity of the Gaia BH3 companion
We have demonstrated that the surface composition of the Gaia BH3 companion may have changed from its intrinsic composition due to metal pollution from the ISM, especially during its main sequence evolution. As the companion evolved beyond the main sequence, it would have experienced mixing as part of its evolutionary process (A. I. Karakas & J. C. Lattanzio, 2014). To quantify the observational signature of external pollution, we utilize MESA to calculate the mass of a star’s convective region as a function of its age, which we characterize using its photospheric radius. This allows us to calculate the surface abundance of a 0.8 $M_{\odot}$ star throughout its evolutionary history for a range of initial birth metallicities. We plot these results in Figure 6.
The diagonal line in Figure 6 indicates when the surface abundance of the star is equal to its birth metallicity. The colored region denotes the parameter space in which the surface metallicity is higher than the birth metallicity due to accretion from the ISM material, as detailed in Figure 4. The darker colors in the diagram represent stars on the main sequence, which have shallower convective envelopes. In contrast, the lighter colors indicate stars in the giant branch, characterized by progressively larger convective envelopes. The orange curve denotes stars on the main sequence. The evolution of the star past the main sequence is then represented as a straight line downwards, expanding in size while decreasing in observed [Fe/H]. The horizontal teal dashed line represents the observed value of [Fe/H] for the Gaia BH3 star and the star-shaped symbol correlates this value with the observed radius of the star. As a greater percentage of the star becomes convective, the metals accumulated over time are mixed more efficiently and become diluted. This dilution tends to diminish any signs of metal pollution, leading to the value of the observed metallicity to approach the value at birth. As a comparison point, we plot a green horizontal line for the case of a star made up solely of hydrogen at birth, accreting metals from the ISM. We argue that this would be the minimum observed metallicity for an $0.8\,M_{\odot}$ accreting from the wake of a $32.7\,M_{\odot}$ companion. For guidance, we also provide the critical metallicity estimates for Population II star formation, $\rm{Z}_{\rm{crit}}$, as derived by V. Bromm & A. Loeb (2003) and R. Schneider et al. (2006), represented here as the two vertical dotted lines.
@@ 5.4 Gaia BH3-like binaries in Globular Clusters
Numerous stellar-mass black hole candidates in binary systems have been successfully identified within Galactic globular clusters (S. R. Kulkarni et al., 1993; J. Strader et al., 2012; A. Bahramian et al., 2017; B. Giesers et al., 2018; T. Shenar et al., 2022). These discoveries are particularly exciting given the significant advancements in the field of gravitational waves in recent years, which suggest that dynamic interactions in dense star clusters could lead to black hole binary mergers (J. Samsing et al., 2014; C. L. Rodriguez et al., 2016; N. C. Weatherford et al., 2018; K. Kremer et al., 2020; C. S. Ye et al., 2024).
Using a typical initial mass function, a globular cluster containing a million initial stars could produce thousands of black holes, many of which are expected to be in binary systems (e.g., A. Vigna-Gómez & E. Ramirez-Ruiz, 2023). ”If a black hole–hosting binary receives a natal kick exceeding the stellar cluster’s escape velocity, it will be ejected, and its subsequent mass accretion history will follow the scenario described in Section 5.1. On the other hand, if they receive a natal kick smaller than the cluster’s escape velocity, they will be retained. In this case, the mass accretion history may be further enhanced by gas retention and prolonged enrichment from evolved populations of cluster stars (N. Prantzos et al., 2007; P. Ventura & F. D’Antona, 2008a; J. Pflamm-Altenburg & P. Kroupa, 2009; C. Conroy et al., 2011; J. P. Naiman et al., 2018a).
In this context, Gaia BH3, linked to the ED-2 retrograde halo stellar stream, is believed to have originated from a disrupted star cluster (E. Balbinot et al., 2024; D. Mar´ın Pina et al., 2024). As a result, it may have undergone additional mass accretion from the retained ejecta of asymptotic giant branch stars within the cluster (J. P. Naiman et al., 2018a). This process could have altered the companion’s abundance patterns for elements such as helium (He), magnesium (Mg), carbon (C), nitrogen (N), oxygen (O), sodium (Na), and aluminum (Al) (e.g., P. Ventura & F. D’Antona, 2008b). The accretion history of stars with black hole companions in stellar clusters is certainly more complex than described in Section 5.1, and its associated chemical signatures could provide further insights into their origins. Future research will assess the long-term impact of accretion on binary systems in globular clusters.
@@ 6 Summary and Conclusions
  • We use hydrodynamical simulations to study the mass accretion rate around the sink particles and the density structures that develop in the vicinity of the binary system (Figure 3). This allows us to estimate the average local density surrounding the less massive stellar companion, which is used to estimate its mass accretion rate. By following this procedure, we derive an expression for how the mass accretion rate onto the less massive companion varies with the mass ratio and the separation of the binary components. Specifically we find that $\dot{M}\propto(M_{\rm{BH}}/M_{*})(R_{a}/a)$.
@@
Our views on the topics discussed herein have been clarified through insightful exchanges with K. El-Badry, R. Foley, B. Gaensler, and E. Kirby. We acknowledge use of the lux supercomputer at UC Santa Cruz, funded by NSF MRI grant AST 1828315. We acknowledge support by the Heising-Simons Foundation and the NSF (AST-2307710, AST-2206243, AST-1911206, and AST-1852393). The 3D hydrodynamics software used in this work was developed in part by the DOE NNSA- and DOE Office of Science-supported FLASH Center for Computational Science at the University of Chicago and the University of Rochester. $S$ acknowledges support from the European High Performance Computing Joint Undertaking (EuroHPC JU) and the Research Council of Norway through the funding of the SPACE Center of Excellence (grant agreement N0101093441). MSF gratefully acknowledges support provided by the Office of the Vice Chancellor for Research and Graduate Education at the University of Wisconsin–Madison with funding from the Wisconsin Alumni Research Foundation.
@@ Appendix B Density redistribution based on adiabatic index
In dense environments, shortly after a binary system forms, cooling is anticipated to play a significant role (e.g., N. Kaaz et al., 2019, 2021). In these situations, the material becomes more compressible, allowing for more effective accretion. Consequently, the gas can be regarded as nearly isothermal rather than adiabatic. To approximate the effects of the radiative cooling on the gas, we vary the adiabatic constant. For the nearly isothermal case, we use $\gamma=1.1$. We also explore $\gamma=4/3$ for completeness.

### v3

@@ 5.2 Altering the birth metallicity of main-sequence companions to LIGO-like black holes
Here we relax the constraints derived from the specific configuration of the BH3 system and expand our calculations to consider metal pollution in a $0.8\,M_{\odot}$ main-sequence star in a binary system with varying separations ($a$ from $0.01$ to $1R_{a}$) and black hole companion masses ($M_{\rm BH}$ from $10$ to $100\,M_{\odot}$). We present the results from this exercise in Figure 5, for which we have derived the total amount of iron accreted using the median value calculated from the particle trajectories of all halo stars (bottom panel) and Gaia BH3–like stars (top panel) for redshifts where $a/R_{a}<1$. We give, for simplicity, an initial birth [Fe/H] value of 5-5 to all stellar companions.
@@
This exercise allows us to probe a diverse collection of binary black hole-star systems that can inform future detections in the Milky Way. The range of stellar black hole masses in the Galaxy is closely linked to the life cycle of massive stars (e.g., 2010ApJ...725.1918O), as well as the energy and dynamics associated with supernova explosions (e.g., 2016ApJ...821...38S; 2018ApJ...862L...3S; 2024PhRvL.132s1403V). By studying this distribution through observations (e.g., 2020A&A...638A..94O), we can gain insights into the types of black hole companions we might expect to find alongside low-mass stars. In the distribution at the top of Figure 5, we plot the mass distribution of the LIGO-Virgo-Kagra black holes (LIGO) and the black hole components of low mass X-ray binaries (Fortin_2024). This distribution takes into account the masses of the primary and secondary components of the mergers, while black holes resulting from the merger were not considered. The low mass X-ray binary distribution peaks at around $15\,M_{\odot}$ while the LIGO distribution peaks at a larger value around $35\,M_{\odot}$.
@@
A few important points stand out from a close examination of Figure 5. For binaries hosting lighter black hole companions at large separations, the resulting surface abundances are low and closest to the star’s birth metallicity. For binaries hosting high-mass black hole companions at short separations, the surface abundance resulting from metal accretion can be significant. On the rightmost panel of Figure 5 we plot, for comparison, the metallicity distribution of the thick disc and halo stars from Gaia DR3 (Kim_2021). This bimodal distribution has a peak near 0.6-0.6, which is attributed to the thick disk stars, and a peak around 1.8-1.8 attributed to the halo stars. We thus conclude that accretion onto low-mass, main-sequence companions, when enhanced by a massive black hole companion, can play a prominent role in altering their surface abundances. In these cases, we expect the inferred surface abundances to trend with stellar luminosity, as mixing processes become progressively more dominant as the star evolves.
@@
Figure 6: The surface abundance enhancement via metal accretion of $0.8\,M_{\odot}$ companion stars with a given birth metallicity. The diagonal gray line shows the values for which the surface abundance is that of the star at birth. The color gradient indicates the different radii of the star as it evolves off the main sequence, with darker colors corresponding to more compact main sequence stars and lighter colors denoting extended giants. The orange curve denotes the surface abundance of stars in the main sequence. The horizontal teal dashed line shows the observed [Fe/H] value for Gaia BH3, and the teal star marker matches the value of the observed metallicity to its current radii. The horizontal green line shows the minimum value for [Fe/H] if, at birth, the star was solely made up of hydrogen, all of its iron content came from accretion from the ISM and the polluting material was mixed into a fully convective star. The dotted vertical lines are critical metallicity for population II formation as derived by Bromm_2003 and Schneider_2006.
@@ 5.3 On the birth metallicity of the Gaia BH3 companion
We have demonstrated that the surface composition of the Gaia BH3 companion may have changed from its intrinsic composition due to metal pollution from the ISM, especially during its main sequence evolution. As the companion evolved beyond the main sequence, it would have experienced mixing as part of its evolutionary process (2014PASA...31...30K). To quantify the observational signature of external pollution, we utilize MESA to calculate the mass of a star’s convective region as a function of its age, which we characterize using its photospheric radius. This allows us to calculate the surface abundance of a 0.8 $M_{\odot}$ star throughout its evolutionary history for a range of initial birth metallicities. We plot these results in Figure 6.
The diagonal line in Figure 6 indicates when the surface abundance of the star is equal to its birth metallicity. The colored region denotes the parameter space in which the surface metallicity is higher than the birth metallicity due to accretion from the ISM material, as detailed in Figure 4. The darker colors in the diagram represent stars on the main sequence, which have shallower convective envelopes. In contrast, the lighter colors indicate stars in the giant branch, characterized by progressively larger convective envelopes. The orange curve denotes stars on the main sequence. The evolution of the star past the main sequence is then represented as a straight line downwards, expanding in size while decreasing in observed [Fe/H]. The horizontal teal dashed line represents the observed value of [Fe/H] for the Gaia BH3 star and the star-shaped symbol correlates this value with the observed radius of the star.
As a greater percentage of the star becomes convective, the metals accumulated over time are mixed more efficiently and become diluted. This dilution tends to diminish any signs of metal pollution, leading to the value of the observed metallicity to approach the value at birth. As a comparison point, we calculate the observed metallicity of a fully convective $0.8\,M_{\odot}$ star made up of solely hydrogen that has been accreting from the wake of a $32.7\,M_{\odot}$ companion for its entire history. We argue that, although unrealistic, this scenario presents the limit of the minimum observed metallicity given than a $100\%$ hydrogen star is the extreme limit of metallicity and a fully convective star is the case of maximal mixing of the accreted material. We plot the value obtained from this calculation as a green horizontal line in Figure 6. For guidance, we also provide the critical metallicity estimates for Population II star formation, $\rm{Z}_{\rm{crit}}$, as derived by Bromm_2003 and Schneider_2006, represented here as the two vertical dotted lines.
@@ 5.4 Gaia BH3-like binaries in Globular Clusters
Numerous stellar-mass black hole candidates in binary systems have been successfully identified within Galactic globular clusters (1993Natur.364..421K; 2012Natur.490...71S; 2017MNRAS.467.2199B; 2018MNRAS.475L..15G; 2022NatAs...6.1085S). These discoveries are particularly exciting given the significant advancements in the field of gravitational waves in recent years, which suggest that dynamic interactions in dense star clusters could lead to black hole binary mergers (2014ApJ...784...71S; 2016PhRvD..93h4029R; 2018ApJ...864...13W; 2020ApJ...903...45K; 2024ApJ...975...77Y).
Using a typical initial mass function, a globular cluster containing a million initial stars could produce thousands of black holes, many of which are expected to be in binary systems (e.g., 2023ApJ...946L...2V). If a black hole–hosting binary receives a natal kick exceeding the stellar cluster’s escape velocity, it will be ejected, and its subsequent mass accretion history will follow the scenario described in Section 5.1. On the other hand, if they receive a natal kick smaller than the cluster’s escape velocity, they will be retained. In this case, the mass accretion history may be further enhanced by gas retention and prolonged enrichment from evolved populations of cluster stars (2007A&A...470..179P; 2008MNRAS.385.2034V; 2009MNRAS.397..488P; 2011ApJ...741...72C; 2018MNRAS.478.2794N).
In this context, Gaia BH3, linked to the ED-2 retrograde halo stellar stream, is believed to have originated from a disrupted star cluster (2024A&A...687L...3B; 2024A&A...688L...2M). As a result, it may have undergone additional mass accretion from the retained ejecta of asymptotic giant branch stars within the cluster (2018MNRAS.478.2794N). This process could have altered the companion’s abundance patterns for elements such as helium (He), magnesium (Mg), carbon (C), nitrogen (N), oxygen (O), sodium (Na), and aluminum (Al) (e.g., 2008A&A...479..805V). The accretion history of stars with black hole companions in stellar clusters is certainly more complex than described in Section 5.1, and its associated chemical signatures could provide further insights into their origins. Future research will assess the long-term impact of accretion on binary systems in globular clusters.
@@ 6 Summary and Conclusions
  • We use hydrodynamical simulations to study the mass accretion rate around the sink particles and the density structures that develop in the vicinity of the binary system (Figure 3). This allows us to estimate the average local density surrounding the less massive stellar companion, which is used to estimate its mass accretion rate. By following this procedure, we derive an expression for how the mass accretion rate onto the less massive companion varies with the mass ratio and the separation of the binary components. Specifically we find that $\dot{M_{*,\rm{bin}}}\propto(M_{\rm{BH}}/M_{*})(R_{a}/a)$.
@@
We thank the referee for their insightful comments, which have significantly improved the clarity of this paper. Our views on the topics discussed herein have been clarified through insightful exchanges with K. El-Badry, R. Foley, B. Gaensler, and E. Kirby. We acknowledge use of the lux supercomputer at UC Santa Cruz, funded by NSF MRI grant AST 1828315. We acknowledge support by the Heising-Simons Foundation and the NSF (AST-2307710, AST-2206243, AST-1911206, and AST-1852393). The 3D hydrodynamics software used in this work was developed in part by the DOE NNSA- and DOE Office of Science-supported FLASH Center for Computational Science at the University of Chicago and the University of Rochester. $S$ acknowledges support from the European High Performance Computing Joint Undertaking (EuroHPC JU) and the Research Council of Norway through the funding of the SPACE Center of Excellence (grant agreement N0101093441). MSF gratefully acknowledges support provided by the Office of the Vice Chancellor for Research and Graduate Education at the University of Wisconsin–Madison with funding from the Wisconsin Alumni Research Foundation.
@@ Appendix B Density redistribution based on adiabatic index
In dense environments, shortly after a binary system forms, cooling is anticipated to play a significant role (e.g., 2019ApJ...876..142K; 2021ApJ...917...36K). In these situations, the material becomes more compressible, allowing for more effective accretion. Consequently, the gas can be regarded as nearly isothermal rather than adiabatic. To approximate the effects of the radiative cooling on the gas, we vary the adiabatic constant. For the nearly isothermal case, we use $\gamma=1.1$. We also explore $\gamma=4/3$ for completeness.
@@
Appendix C Black hole feedback
One of the main uncertainties in the accretion flow surrounding the companion is the role of feedback from the BH companion. Feedback can be deposited into the gas either radiatively or mechanically, through winds, and can affect the mass accretion rate onto the stellar companion. At the low mass accretion rates relevant to our system, accretion is expected to occur radiatively inefficiently. In such cases, known as the adiabatic inflow–outflow solution (ADIOS; 1999MNRAS.303L...1B; 2004MNRAS.349...68B), accretion tends to be strongly inhibited, with much less gas reaching the black hole than is supplied at large distances. This is because the outward transport of angular momentum also involves the transport of energy, which can cause gas far from the black hole to become unbound.
For perfectly adiabatic (non-radiating) systems, the mass inflow rate depends on radius as $\dot{M}_{w}\propto R$, while the wind kinetic luminosity scales as $L_{w}\propto R^{1/2}$. This implies that most of the wind’s kinetic energy is generated at large radii (2012MNRAS.420.2912B):
$\begin{split}L_{w}&\approx\dot{M}_{w}v_{w}^{2}\\ &\approx\eta\dot{M}_{\rm{B}}v_{w}^{2},\end{split}$ (C1)
where $\dot{M}_{w}$ is the mass loss from the black hole due to winds and $v_{w}$ is the velocity of the ejected wind. Here $\eta=R_{\rm{disk}}/R_{a}$ and $R_{\rm{disk}}$ is the characteristic radius at which the infalling gas, conserving its angular momentum, begins to form a stable, approximately Keplerian disk.
To determine the value of $R_{\rm{disk}}$, we estimate the circularization radius based on the distance between the black hole and the first Lagrangian point of the binary system, $R_{L1}$. This distance can be approximated as
$R_{L1}\approx a\left(0.5+0.277\log q\right).$ (C2)
The disk radius is then defined as,
$R_{\rm{disk}}\approx a(1+q)\left(\frac{R_{L1}}{a}\right)^{4}.$ (C3)
Using this definition, we can find the value of $\eta$,
$\eta={R_{\rm disk}\over R_{a}}\approx\frac{a}{R_{a}}(1+q)\left(0.5+0.277\log q\right)^{4}.$ (C4)
For the Gaia BH3 system, $q=1/40$ and $a/R_{a}$ varies with time. The value of $\eta$ will be greatest at larger separations. We therefore take $a/R_{a}=1$ to represent the value with the greatest disk extension that still falls within the regime of binary accretion. Using this value and the mass ratio of approximately 40 for Gaia BH3, we find $\eta=3.5\times 10^{-4}$.
If an accreting BH produces a wind, it can potentially carve out a cavity in the surrounding gas, reducing or even suppressing the gas supply available for accretion. We calculate the shock radius of this cavity, $R_{s}$, as the location where the ram pressure of the wind balances the ambient gas pressure, which can be expressed as
$R_{s}=\sqrt{\frac{\dot{M}_{w}v_{w}}{4\pi\rho_{\infty}v_{\infty}^{2}}}.$ (C5)
To compare this to the relevant physical scale, we can express the shock radius in terms of the Bondi radius,
$R_{a}=\sqrt{\frac{\dot{M}_{\rm{B}}}{\pi\rho_{\infty}v_{\infty}}}.$ (C6)
This ratio can then be written as
$\begin{split}\frac{R_{s}}{R_{a}}&=\sqrt{\frac{\dot{M}_{w}v_{w}}{4\dot{M}_{\rm{B}}v_{\infty}}}\\ &=\frac{1}{2}\sqrt{\frac{\dot{M}_{w}}{\dot{M}_{\rm{B}}}}\sqrt{\frac{v_{w}}{v_{\infty}}},\end{split}$ (C7)
where $\dot{M}_{w}=\eta\dot{M}_{\rm B}$ and $\frac{v_{w}}{v_{\infty}}\approx\eta^{-1/2}$, so that
$\frac{R_{s}}{R_{a}}\approx\frac{1}{2}\eta^{1/4},$ (C8)
Here, $\eta\ll 1$. This suggests that for radiatively inefficient accretion flows, the luminosity produced by the black hole is unlikely to carve out a large cavity or significantly impede the gas supply onto the companion. Nevertheless, the flow may still be reshaped, potentially affecting the scaling laws derived in this paper (Section 4), which were obtained under the assumption of no feedback.